OpenStudy (anonymous):

What is the least common multiple of x squared+8x+15 and x squared + 5x +6?

OpenStudy (anonymous):

factor them first = (x+5)(x+3) second = (x+3)(x+2) answer = (x+3)

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

@elecengineer: He's lookign for the least common multiple.

OpenStudy (radar):

I believe elecengineer gave him a start, couldn't he just multiply (x=5)(x+3)(x=2) and come up with the LCM?

OpenStudy (radar):

(x+5)(x+3)(x+2) I mean

OpenStudy (anonymous):

Yeah. The LCM=(x+5)(x+3)(x+2).

OpenStudy (radar):

There you go barrelracer38. Been to a rodeo lately??lol

OpenStudy (anonymous):

LOL